# Navigating High Dimensional Concept Space with Metalearning

### v1

@@ Abstract
Rapidly learning abstract concepts from limited examples is a hallmark of human intelligence. This work investigates whether gradient-based meta-learning can equip neural networks with inductive biases for efficient few-shot acquisition of discrete concepts. We compare meta-learning methods against a supervised learning baseline on Boolean tasks generated by a probabilistic context-free grammar (PCFG). By systematically varying concept dimensionality (number of features) and compositionality (depth of grammar recursion), we identify regimes in which meta-learning robustly improves few-shot concept learning. We are able to show that meta-learners are much better able to handle compositional complexity than featural complexity and establish an empirical analysis demonstrating how featural complexity shapes ’concept basins’ of the loss landscape, allowing curvature-aware optimization to be more effective than first order methods. We see that we can robustly increase generalization on complex concepts by increasing the number of adaptation steps in meta-SGD, encouraging exploration of rougher loss basins. Overall, this work highlights the intricacies of learning compositional versus featural complexity in high dimensional concept spaces and provides a road to understanding the role of 2nd order methods and extended gradient adaptation in few-shot concept learning.
@@ 1 Introduction
Human learners are able to learn concepts from remarkably little data [10]. Early models of concept learning that were able to match human performance leveraged Bayesian inference and symbolic modeling [5, 15, 3], explaining human performance as an approximation of rational inference. However, these methods often suffered from computational intractability. Recent approaches have combined Bayesian modeling with neural networks through metalearning, allowing models to match Bayesian performance with a fraction of the data [9], thus providing a more plausible mechanism to explain the remarkable sample efficiency of human concept learning. Few approaches have tested the opposite end of the spectrum and explored entirely gradient-based meta-learning methods and effects on sample efficiency and generalization in the concept learning domain. Furthermore, while meta-learning has achieved impressive results across domains including perception, control, and reasoning, existing evaluations often focus on performance within fixed datasets, leaving underexplored how meta-learning behaves as task complexity systematically increases to higher dimensionalities.
@@
In this work, we study the limits of end-to-end, gradient-based meta-learning in a Boolean concept learning task, in a setting that allows precise control over compositional and featural complexity via a probabilistic context free grammar (PCFG). By extending a PCFG-based concept generator developed initially by Goodman et al. 2008 [5], we independently vary featural dimensionality (number of binary input features) and compositional depth (logical recursion) to create a curriculum of tasks with increasing structural complexity. This setting enables us to ask how the effectiveness of meta-learning scales with the complexity of the underlying concept space.
@@
We compare gradient-based meta-learning (Meta-SGD) against standard supervised learning (SGD) on few-shot Boolean classification tasks. Results show that meta-learning is incredibly robust at handling increased compositional depth in line with previous work [2, 9], but suffers with increased featural dimensionality, mirroring limitations on human working memory capacity [12]. To explain this, we show how increasing dimensionality results in an increase in the roughness of the loss landscape of the ’concept basin’. As a potential remedy, we find empirical evidence that increasing the number of adaptation steps can reliably help a meta-learner navigate these rougher loss landscapes, to find generalization-friendly weight initializations. We also present early evidence of how curvature awareness (the second-order gradient term) helps meta-learning effectively navigate different concept complexities by analyzing the Hessian trace of increasingly difficult concept losses. An analysis of loss landscape roughness reveals a strong correlation between landscape curvature and relative gains from meta-learning, proposing a mechanistic account of when and why meta-learning is effective in higher dimensions.
@@ 2 Related Work
Compositional generalization and concept learning. Symbolic rule induction methods, such as Bayesian Program Learning (BPL) [10] and the Rational Rules model [5], achieve human-level one-shot learning by leveraging explicit grammars. However, they require handcrafted generative models and search. Neural sequence-to-sequence models struggle with systematic generalization on tasks like SCAN [8], and neural meta-learners underperform on benchmarks like CURI [16]. Meta-learning has recently been used to improve compositional generalization in NLP [2] and neuro-symbolic reasoning systems [17], but its role in Boolean concept induction remains underexplored. A theoretical framework for compositional generalization in neural networks was recently proposed [1], and surveys highlight the challenges and opportunities for compositional AI [14]. I study this in a controlled discrete (Boolean) setting to isolate logical structure.
@@ 3 Experimental Setup
All methods use a 5-layer MLP (128 hidden units/layer, ReLU, sigmoid output). We compare models trained with three stochastic gradient descent (SGD) learning algorithms, varying the order of the gradients and adaptation steps: 1st-Order and 2nd-Order Meta-SGD with both 1 and 10 adaptation (gradient) steps and regular SGD: training from scratch per task using Adam (learning rate 0.001) on $S_{C}$.
@@
Increasing $K_{adapt}$ allows more extensive search in the task-specific loss landscape, incrementally adjusting the MLP’s decision boundaries to correctly classify support set examples. Meta-SGD models were meta-trained for 10,000 episodes. All evaluations were averaged over 5 random seeds on 1,000 unseen tasks (like those shown in Figure 1). For trajectory comparisons, SGD is trained for steps equivalent to processing a fixed total number of samples. Performance is assessed using final mean accuracy (Appendix A.3) and data efficiency (samples required to reach 60% accuracy, Appendix A.1).
@@ 5 Loss Landscape Analysis
To explain meta-learning’s successes and failures further, we analyzed loss landscape topology across concept complexities, revealing causal connections between landscape properties and meta-learning performance.
@@ 5.1 Methodology
We define roughness as optimization instability: trajectory variation during training. Our metric extracts loss sequences $L=[l_{1},l_{2},\ldots,l_{T}]$, normalizes to 200 episodes, applies Gaussian smoothing ($\sigma=1$), computes discrete second derivatives $\nabla^{2}L_{i}=l_{i+1}-2l_{i}+l_{i-1}$, and calculates:
@@
This enables quantitative relationships between Boolean concept complexity, landscape roughness, and meta-learning effectiveness.
@@ 5.2 Complexity-Dependent Landscape Topology
We analyzed loss landscapes by sampling random directions in parameter space. Boolean concept complexity fundamentally determines landscape topology, creating predictable optimization challenges (Figure 3).
Figure 3: Meta-learning and SGD operate on the same concept loss landscapes (determined by task structure and architecture), but meta-learning learns more efficient navigation strategies (shorter paths to solution point). Top row: 2D loss landscapes for simple, medium, and complex Boolean concepts show identical topology regardless of optimization method. Middle row: 3D visualizations reveal the terrain both algorithms must navigate, with complexity-dependent ruggedness. Note: loss landscapes are local approximations of the full loss landscape.
Our quantitative analysis reveals systematic patterns: simple concepts (2-3 literals) exhibit smooth, quasi-convex landscapes with few local minima (0.3±0.1, roughness = 0.0002); medium concepts (4-6 literals) show moderately rugged topology (1.2±0.4 minima,  4x roughness increase); complex concepts (7+ literals) display highly rugged landscapes with multiple local minima (2.8±0.6,  12x roughness increase).
Thus, featural complexity creates characteristic landscape patterns, which maps to optimization difficulty through the number and distribution of local minima.
@@ 6 Discussion
Compositional Complexity: Meta-Learning’s Strength. Across all experiments, meta-learning demonstrates remarkable robustness to increasing compositional depth. Moving from $D=3$ to $D=7$ (simple to deeply nested logical structures) shows minimal performance degradation for Meta-SGD, while SGD suffers substantially. Our loss landscape analysis reveals why: compositional complexity primarily affects the logical structure within concept space but preserves relatively navigable optimization surfaces. The PCFG’s recursive depth creates more intricate Boolean relationships without fundamentally altering the smoothness of parameter space traversal. Meta-learning’s learned initialization and adaptive step sizes prove particularly effective at discovering these hierarchical patterns within reasonable adaptation budgets.
Featural Complexity: The Fundamental Challenge. In stark contrast, increasing featural dimensionality poses severe challenges for all methods, with performance collapsing dramatically at $F=32$. This reveals a deeper truth about the nature of concept learning: while logical complexity (compositionality) can be handled through better optimization strategies, dimensional complexity fundamentally alters the search space structure. The explosion from $2^{8}$ to $2^{32}$ possible input configurations under high data sparsity creates loss landscapes so rugged and high-dimensional that meta-learning alone cannot overcome the curse of dimensionality. Our roughness analysis confirms that featural complexity creates exponentially more challenging optimization terrain than compositional complexity and provides insight for future concept learning work.
Landscape Implications. This dual-axis analysis reveals that not all forms of ”complexity” are equivalent from an optimization perspective. Compositional depth affects the logical relationships that must be learned but preserves loss surface properties. The extent to which this is true in higher dimensional setting and with more complex models deserves further investigation. Featural dimensionality, however, fundamentally transforms the geometry of the optimization problem itself.
These findings suggest that meta-learning is particularly well-suited for domains where complexity arises from structural relationships rather than raw dimensionality, explaining its success in few-shot learning across compositionally rich but feature-moderate domains [13, 6].
@@ 7 Future Work
While we allude to similar findings in memory capacity, sample efficiency, and generalization in human learning, this work would benefit from a direct comparison to human performance. The effect of blocking and spacing over interspersed intervals, as done by Goodman et al. (2008) would also be a fruitful area of insight into the dynamics of meta-learning. Finally, concept length generalization to out-of-distribution grammars would allow us to test the effectiveness of the inductive bias given by metalearning.
@@ 8 Conclusion
Our investigation across featural dimensionality and compositional depth reveals when and why meta-learning succeeds in concept acquisition, demonstrating that different complexity types pose different optimization challenges.
Meta-learning exhibits asymmetric robustness across complexity dimensions. While compositional complexity (increasing logical depth from $D=3$ to $D=7$) poses minimal challenges for Meta-SGD, featural complexity (expanding from $F=8$ to $F=32$ features) creates optimization difficulties. Our loss landscape analysis highlights a potential explanation: compositional depth affects logical structure while preserving navigable parameter spaces, whereas featural dimensionality fundamentally transforms optimization geometry.
@@ A.1 Data Efficiency
Figure 4: Training samples to reach 60% validation accuracy (log scale). Meta-SGD methods demonstrate substantial sample efficiency advantages over SGD, particularly in higher depth settings. The 32-feature case only saw first order Meta-SGD with k=10 generalize to above 60 percent accuracy.
@@ A.2 Adaptation Steps Scale with Landscape Complexity
A tricky question in meta-learning is how many gradient steps to use during adaptation. Our analysis reveals that the benefit of additional gradient steps (K=10 vs K=1) scales directly with concept complexity. Figure 5 demonstrates this linear relationship across the spectrum of concept categories we tested above.
@@
Simple concepts show modest 5-8% accuracy improvement from K=10 over K=1 (efficiency ratio 1.4×), medium concepts show substantial 10-12% improvement (efficiency ratio 1.8×), and complex concepts show large 15-20% improvement (efficiency ratio 2.5×). Simple concepts have smooth landscapes navigable with single adaptation steps, while complex concepts have rugged landscapes requiring multiple steps to escape local minima and find better solutions.
A.3 Final Mean Validation Accuracies
Figure 6: Final mean validation accuracies across the matrix of feature dimensionalities (rows/facets) and concept depths (x-axis).
A.4 Curvature Analysis
We compute four curvature-related metrics to characterize landscape geometry: roughness (variance of loss gradients along random directions), Hessian trace ($\text{tr}(\mathbf{H})=\sum_{i}\lambda_{i}$ indicating local curvature), spectral norm ($\|\mathbf{H}\|_{2}=\max_{i}|\lambda_{i}|$ measuring maximum curvature), and condition number ($\kappa(\mathbf{H})=\lambda_{\max}/\lambda_{\min}$ quantifying eigenvalue ratios).
We observe more systematic patterns in how curvature properties scale with concept complexities. For simple concepts (F8D3), Meta-SGD reduces the Hessian trace by 92.6% compared to SGD ($\Delta\text{tr}(\mathbf{H})=-0.926$). Medium concepts (F8D5) show 95.8% trace reduction with 50.9% roughness improvement ($\Delta\text{tr}(\mathbf{H})=-0.958$, $\Delta\sigma^{2}_{\nabla}=-0.509$). Complex concepts (F32D3) maintain 88.5% trace reduction despite landscape complexity ($\Delta\text{tr}(\mathbf{H})=-0.885$).
This is a first step towards beginning to explain how meta-SGD finds better solutions by smoothening the geometry of the optimization trajectory (Figure 3), enabling efficient few-shot learning. This connects to recent theoretical work on meta-learning optimization landscapes [6, 13]. Our findings suggest that meta-learning’s effectiveness stems from its ability to reduce local curvature (creating smoother gradient flows), improve conditioning (reducing eigenvalue ratios $\kappa(\mathbf{H})$ for better convergence), and minimize roughness (eliminating sharp local minima that trap gradient descent). This geometric perspective offers a new lens for understanding meta-learning: rather than simply providing better initializations, meta-learning algorithms actively reshape the optimization trajectory to enable efficient navigation and adaptation.
A.4.1 Relationship to Boolean Concept Structure
The discrete nature of Boolean concepts creates characteristic patterns in loss landscape topology. Our analysis reveals that logical complexity directly correlates with landscape roughness, PCFG depth influences the number and distribution of local minima, and feature dimensionality affects the overall scale of curvature variations.
These findings establish a bridge between symbolic concept structure and continuous optimization geometry, providing theoretical foundation for understanding why meta-learning excels at Boolean concept learning tasks.

### v3

# Adapting to High Dimensional Concept Space with Metalearning

@@ Abstract
Rapidly learning abstract concepts from limited examples is a hallmark of human intelligence. This work investigates whether gradient-based meta-learning can equip neural networks with inductive biases for efficient few-shot acquisition of discrete concepts. I compare meta-learning methods against a supervised learning baseline on Boolean concepts (logical statements) generated by a probabilistic context-free grammar (PCFG). By systematically varying concept dimensionality (number of features) and recursive compositionality (depth of grammar recursion), I delineate between complexity regimes in which meta-learning robustly improves few-shot concept learning and regimes in which it does not. Meta-learners are much better able to handle compositional complexity than featural complexity. I highlight some reasons for this with a loss landscape analysis demonstrating how featural complexity increases the roughness of loss trajectories, allowing curvature-aware optimization to be more effective than first-order methods. I find improvements in out-of-distribution generalization on complex concepts by increasing the number of adaptation steps in meta-SGD, where adaptation acts as a way of encouraging exploration of rougher loss basins. Overall, this work highlights the intricacies of learning compositional versus featural complexity in high dimensional concept spaces and provides a road to understanding the role of 2nd order methods and extended gradient adaptation in meta-learning.
@@ 1 Introduction
Human learners are able to learn concepts from remarkably little data [Lake et al. [2015]]. Early models of concept learning that were able to match human performance leveraged Bayesian inference and symbolic modeling [Goodman et al. [2008], Tenenbaum and Griffiths [2001], Feldman [2000]], explaining human performance as an approximation of rational inference. However, these methods often suffered from computational intractability. Recent approaches have combined Bayesian modeling with neural networks through metalearning, allowing models to match Bayesian performance with a fraction of the data [Lake and Baroni [2023]], thus providing a more plausible mechanism to explain the remarkable sample efficiency of human concept learning. Few approaches have tested the opposite end of the spectrum and explored entirely gradient-based meta-learning methods and effects on sample efficiency and generalization in the concept learning domain. Furthermore, while meta-learning has achieved impressive results across domains including perception, control, and reasoning, existing evaluations often focus on performance within fixed datasets, leaving underexplored how meta-learning behaves as task complexity systematically increases to higher dimensionalities.
@@
In this work, I study the limits of end-to-end, gradient-based meta-learning in a Boolean concept learning task, in a setting that allows precise control over compositional and featural complexity via a probabilistic context free grammar (PCFG). By extending a PCFG-based concept generator developed initially by Goodman et al. 2008 [Goodman et al. [2008]], I independently vary featural dimensionality (number of binary input features) and compositional depth (logical recursion) to create a distribution of tasks with increasing structural complexity. This setting enables a study of how out-of-distribution (OOD) generalization from meta-learning scales with the complexity of the underlying concept space.
@@
I compare gradient-based meta-learning (Meta-SGD) against standard supervised learning (SGD) on few-shot Boolean classification tasks. Results show that meta-learning is incredibly robust at handling increased compositional recursions, in line with previous work [Conklin et al. [2021], Lake and Baroni [2023]], but suffers with increased featural dimensionality. To explain this, I show how increasing dimensionality results in an increase in the roughness of the loss landscape of the ’concept basin’. As a potential remedy, I find empirical evidence that increasing the number of adaptation steps can reliably help a meta-learner navigate these rougher loss landscapes, to find generalization-friendly weight initializations. I also present early evidence of how curvature awareness (the second-order gradient term) helps meta-learning effectively navigate different concept complexities by analyzing the Hessian trace of increasingly difficult concept landscapes. An analysis of loss landscape roughness reveals a strong correlation between landscape curvature and relative gains from meta-learning, partially explaining when and why meta-learning is effective in high dimensional concept learning.
@@ 2 Related Work
Compositional generalization and concept learning. Symbolic rule induction methods, such as Bayesian Program Learning (BPL) [10] and the Rational Rules model [5], achieve human-level one-shot learning by leveraging explicit grammars. However, they require handcrafted generative models and search. Neural sequence-to-sequence models struggle with systematic generalization on tasks like SCAN [8], and neural meta-learners underperform on benchmarks like CURI [15]. Meta-learning has recently been used to improve compositional generalization in NLP [2] and neuro-symbolic reasoning systems [16], but its role in Boolean concept induction remains underexplored. A theoretical framework for compositional generalization in neural networks was recently proposed [1], and surveys highlight the challenges and opportunities for compositional AI [13]. Few studies have explicitly engaged in an exploration of how increasingly recursive structure affects the ability of a meta-learner. In this work, I take a first step in this direction in a discrete, Boolean setting to isolate logical complexity as a predictive variable for metalearning performance.
@@ 3 Experimental Setup
All methods use a 5-layer MLP (128 hidden units/layer, ReLU, sigmoid output). I compare models trained with three stochastic gradient descent (SGD) learning algorithms, varying the order of the gradients and adaptation steps: 1st-Order and 2nd-Order Meta-SGD with both 1 and 10 adaptation (gradient) steps and regular SGD: training from scratch per task using Adam (learning rate 0.001) on $S_{C}$.
@@
Increasing $K_{adapt}$ allows more extensive search in the task-specific loss landscape, incrementally adjusting the MLP’s decision boundaries to correctly classify support set examples. Meta-SGD models were meta-trained for 10,000 episodes. All evaluations were averaged over 5 random seeds on 1,000 unseen tasks (like those shown in Figure 1). For trajectory comparisons, SGD is trained for steps equivalent to processing a fixed total number of samples. Performance is assessed using final mean accuracy and data efficiency (samples required to reach 60% accuracy, Appendix A.1).
@@ 5 Loss Landscape Analysis
Behaviorally, we find that meta-learning is highly effective across concept depth, but offers rapidly diminishing returns as we scale the number of features per concept, correpsonding to the depth and breadth of the PCFG in Figure 1. In this section, I take a closer look at how the learning process diverges between meta-SGD and SGD over the concept classes tested above by visualizing the loss landscapes during the training process itself.
@@ 5.1 Methodology
I define roughness as optimization instability: trajectory variation during training. Equation (1) defines a metric of roughness that extracts loss sequences $L=[l_{1},l_{2},\ldots,l_{T}]$, normalized over 200 episodes of training, applies Gaussian smoothing ($\sigma=1$), computes discrete second derivatives $\nabla^{2}L_{i}=l_{i+1}-2l_{i}+l_{i-1}$, and calculates:
@@
The goal is to measure how introducing meta-learning (2nd order gradients) changes the loss function across increasingly complex concepts, each define by their own loss landscapes (Figure 3).
@@ 5.2 Complexity-Dependent Landscape Topology
I analyzed loss landscapes by sampling random directions in parameter space to approximately visualize the optimization challenge posed by each class of boolean concept difficulty. Boolean concept complexity increases roguhness of landscape topology, creating different optimization challenges (Figure 3).
Figure 3: Meta-SGD and SGD operate on the same concept loss landscapes (determined by task structure and architecture), but meta-learning learns more efficient navigation strategies (shorter paths to the solution point - the bottom-most point in each loss ’basin’). Top row: 2D loss landscapes for simple, medium, and complex Boolean concepts show identical topology regardless of optimization method. Middle row: 3D visualizations reveal the terrain both algorithms must navigate, with complexity-dependent roughness. Note: due to computational intractability, loss landscapes are local approximations.
A quantitative analysis reveals systematic patterns: simple concepts (2-3 literals) exhibit smooth, quasi-convex landscapes with few local minima (on average 0.3±0.1, roughness = 0.0002); medium concepts (4-6 literals) show moderately rugged topology (1.2±0.4 minima,  4x roughness increase); complex concepts (7+ literals) display highly rugged landscapes with multiple local minima (2.8±0.6,  12x roughness increase). In other words, featural complexity creates characteristic landscape patterns, which map to optimization difficulties - the number and distribution of local minima. This partially explains the failures of meta-SGD to learn featurally dense concepts (
@@ 6 Discussion
Compositional Complexity Across all experiments, meta-learning demonstrates remarkable robustness to increasing compositional depth. Moving from $D=3$ to $D=7$ (simple to deeply nested logical structures) shows minimal performance degradation for Meta-SGD, while SGD suffers substantially. Our loss landscape analysis reveals why: compositional complexity primarily affects the logical structure within concept space but preserves relatively navigable optimization surfaces. The PCFG’s recursive depth creates more intricate Boolean relationships without altering the smoothness of parameter space traversal. Meta-learning’s learned initialization and adaptive step sizes prove particularly effective at discovering these hierarchical patterns within reasonable adaptation budgets.
Featural Complexity In stark contrast, increasing featural dimensionality poses severe challenges for all methods, with performance collapsing dramatically at $F=32$. This reveals a deeper truth about the nature of concept learning: while logical complexity (compositionality) can be handled through better optimization strategies, dimensional complexity alters the search space. The explosion from $2^{8}$ to $2^{32}$ possible input configurations under high data sparsity creates loss landscapes so rugged and high-dimensional that meta-learning alone cannot overcome the curse of dimensionality. Our roughness analysis confirms that featural complexity creates exponentially more challenging optimization terrain than compositional complexity and provides insight for future concept learning work.
Landscape Implications. This dual-axis analysis reveals that not all forms of ”complexity” are equivalent from an optimization perspective. Compositional depth affects the logical relationships that must be learned but preserves loss surface properties. The extent to which this is true in higher dimensional setting and with more complex models deserves further investigation. Furthermore, these loss trajectories were randomly sampled in a small, locally convex area of the total loss landscape, but further probing with more compute could reveal more complex topologies than the approximations provided.
These findings suggest that meta-learning is particularly well-suited for domains where complexity arises from structural relationships rather than raw dimensionality, explaining its success in few-shot learning across compositionally rich but feature-constrained domains [12, 6].
@@ 7 Future Work
An unanswered and intriguing question in the study of compositionality is how compositional concepts are embedded in the high-dimensional vector space parametrized by the hidden units of a sufficiently deep neural network. In this case, we have already seen that simple MLP’s can accurately classify boolean concepts of low to medium compositional depth. I would be very curious to examine the learned representations of the MLP’s to explore if any regular structure emerges. Given that the concept space itself is regular and well-defined hierarchically, I would suggest a simple Principal Component Analysis (PCA) on the hidden units of the MLP’s trained in the above experiments. In a separate vein of inquiry, the extent to which I tested out of distrbution grammars was very limited - concept length generalization to out-of-distribution grammars would allow us to test the effectiveness and range of the inductive bias endowed by metalearning.
@@ 8 Conclusion
This investigation across featural dimensionality and compositional depth reveals when and why meta-learning succeeds in boolean concept acquisition.
Meta-learning exhibits asymmetric robustness across complexity dimensions. While compositional complexity poses minimal challenges for Meta-SGD, featural complexity creates challenging optimization problems. A loss landscape analysis highlights a potential explanation: compositional depth affects logical structure while preserving navigable parameter spaces, whereas featural dimensionality transforms loss landscapes, creating ”rougher” basins in which second-order methods become relatively more effective than first-order methods.
@@ A.1 Data Efficiency
Figure 4: Training samples to reach 60% validation accuracy (log scale). While results are mixed and vanilla SGD never achieves the floor accuracy of 60%, there are intriguing early trends for first versus second order Meta-SGD. FOr example, while first order Meta-SGD outperforms in high depth settings with a low number of features, second order greatly outperforms when increasing featural complexity. The 32-feature case (bottom panel) has only data for first order Meta-SGD with k=10 steps because it was the only method to generalize to above 60 percent accuracy.
@@ A.2 Adaptation Steps Scale with Landscape Complexity
A tricky question in meta-learning is how many gradient steps to use during test-time adaptation. The previous analysis reveals that the benefit of additional gradient steps (K=10 vs K=1) scales directly with concept complexity. Figure 5 demonstrates this linear relationship across the spectrum of concept categories I tested above.
@@
Simple concepts show modest 5-8% accuracy improvement from K=10 over K=1 (efficiency ratio 1.4×), medium concepts show substantial 10-12% improvement (efficiency ratio 1.8×), and complex concepts show large 15-20% improvement (efficiency ratio 2.5×). From the loss landscape analysis, these adaptation gains make sense: simple concepts have smooth landscapes navigable with single adaptation steps, while complex concepts have rugged landscapes requiring multiple steps to escape local minima and find better solutions.
A.3 Curvature Analysis
I compute four curvature-related metrics to characterize landscape geometry: roughness (variance of loss gradients along random directions), Hessian trace ($\text{tr}(\mathbf{H})=\sum_{i}\lambda_{i}$ indicating local curvature), spectral norm ($\|\mathbf{H}\|_{2}=\max_{i}|\lambda_{i}|$ measuring maximum curvature), and condition number ($\kappa(\mathbf{H})=\lambda_{\max}/\lambda_{\min}$ quantifying eigenvalue ratios).
With these metrics as proxies for optimization difficulty, I find several systematic patterns relating curvature to concept complexity. For simple concepts (e.g. F8D3), Meta-SGD reduces the Hessian trace by 92.6% compared to SGD ($\Delta\text{tr}(\mathbf{H})=-0.926$). Medium concepts (F8D5) show 95.8% trace reduction with 50.9% roughness improvement ($\Delta\text{tr}(\mathbf{H})=-0.958$, $\Delta\sigma^{2}_{\nabla}=-0.509$). Complex concepts (F32D3) maintain 88.5% trace reduction ($\Delta\text{tr}(\mathbf{H})=-0.885$). Hessian trace reduction in this case can be seen as a form of regularization, guiding the networks towards finding flatter minima in the loss landscape, which in turn promote better generalization. Although hard to visualize, we can refer back to Figure 3 to gain a better intuition for how this may work in complex concept space, where the relative change in Hessian trace between sharp and smooth minima is largest.
This is meant to serve as a step towards beginning to explain how meta-SGD finds better solutions by smoothening the geometry of the optimization trajectory (Figure 3), enabling efficient few-shot learning. Recent theoretical work on meta-learning optimization landscapes [6, 12] highlight similar findings. These initial metrics suggest that meta-learning’s effectiveness stems from its ability to reduce local curvature (creating smoother gradient flows), improve conditioning (reducing eigenvalue ratios $\kappa(\mathbf{H})$ for better convergence), and minimize roughness (eliminating sharp local minima that trap gradient descent). This offers a different lens for understanding meta-learning: rather than simply providing better initializations, Meta-SGD actively reshapes the optimization trajectory to enable efficient navigation and adaptation.